Andrew wants to divide his rectangular garden into two parts diagonally. The garden has a length of 40 feet and width 30 feet.
What will be the area of each section of the garden after the division?

Answer by pmesler(52)

First, draw a rectangle. Label the width 30 and the length 40. Then draw a diagonal across the entire rectangle. This forms two right triangles. This diagonal is the hypotenuse that both right triangles share, one for each side. To find the area of a triangle use the formula
A = b * h/2. Here the base, b is the length, and the width w, is the height h.
Therefore the area is
A = 30 * 40 / 2 = 120/2 = 60
Therefore the area for each section of the garden will be 60 square feet.